The goose has a mass of 20.3 lb (pounds) and is flying at 7.50 miles/h (miles per hour). What is the kinetic energy of the goose in joules?
Answer #1

1 lb = 453.59 g = 0.45359 kg

thus

20.3 lb = 20.3 * 0.45359 = 9.208 kg

7.50 miles/h = 7.50 * 1609.34 m / 3600 s = 3.35 m / s

therefore,

kinetic energy = 1 / 2 * m * V^2 = 1 / 2 * 9.208 * (3.35)^2 = 51.75 J

thus

kinetic energy of the goose = 51.75 J